Consider the following simplified APT model: Expected R Premium (%) 8.4 .5 5.5 Factor Market Interest rate Yield spread Factor Risk Exposures Yield Spread (b3) Market Interest Rate (02) Sbock (b) 1.6 1.8 1,8 1.0 Consider a portolio with Oqual investments in stocks P, P, and P, Assurme re- a. What are the factor risk exposures for the portfolio? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to 3 decimal places Factor Risk Exposures Market (D) Interest rate (2 Yield spread b. What is the portfolios expected return? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places) Expected return
Answer #1
a) Factor risk exposure for the portfolio:
Market (b1) = (1.6+1.8+0.3)/3 = 1.233
Interest rate (b2) = (-1.8+0+0.9)/3 = -0.300
Yield spread (b3) = (-0.8+0.6+1.0)/3 = 0.267
b) Portfolio's expected return = 5+1.233*8.4+(-0.300*-0.5)+0.267*5.5) = 16.98%